On October 29, 2016, Lobo Co. began operations by purchasing razors for resale. Lobo uses the perpetual inventory method. The razors have a 90-day warranty that requires the company to replace any nonworking razor. When a razor is returned, the company discards it and mails a new one from Merchandise Inventory to the customer. The company's cost per new razor is $15 and its retail selling price is $80 in both 2016 and 2017. The manufacturer has advised the company to expect warranty costs to equal 7% of dollar sales. The following transactions and events occurred.

2016

Nov.

11

Sold 70 razors for $5,600 cash.

30

Recognized warranty expense related to November sales with an adjusting entry.

Dec.

9

Replaced 14 razors that were returned under the warranty.

16

Sold 210 razors for $16,800 cash.

29

Replaced 28 razors that were returned under the warranty.

31

Recognized warranty expense related to December sales with an adjusting entry.


2017

Jan.

5

Sold 140 razors for $11,200 cash.

17

Replaced 33 razors that were returned under the warranty.

31

Recognized warranty expense related to January sales with an adjusting entry.

1.1 Prepare journal entries to record above transactions and adjustments for 2016.

1.2 Prepare journal entries to record above transactions and adjustments for 2017.

2. How much warranty expense is reported for November 2016 and for December 2016?

3. How much warranty expense is reported for January 2017?

4. What is the balance of the Estimated Warranty Liability account as of December 31, 2016?

5. What is the balance of the Estimated Warranty Liability account as of January 31, 2017?
Answer #1
1.1
Date General Journal Debit Credit
11-Nov Cash 5,600
Sales 5,600
11-Nov Cost of goods sold 1050 =70*15
Merchandise inventory 1,050
30-Nov Warranty expense 392 =5600*7%
Estimated warranty liability 392
9-Dec Estimated warranty liability 210 =14*15
Merchandise inventory 210
16-Dec Cash 16,800
Sales 16,800
16-Dec Cost of goods sold 3,150 =210*15
Merchandise inventory 3,150
29-Dec Estimated warranty liability 420 =28*15
Merchandise inventory 420
31-Dec Warranty expense 1,176 =16800*7%
Estimated warranty liability 1,176
1.2
Date General Journal Debit Credit
5-Jan Cash 11,200
Sales 11,200
5-Jan Cost of goods sold 2,100 =140*15
Merchandise inventory 2,100
17-Jan Estimated warranty liability 495 =33*15
Merchandise inventory 495
31-Jan Warranty expense 784 =11200*7%
Estimated warranty liability 784
2
Warranty expense for November 392 =5600*7%
Warranty expense for December 1176 =16800*7%
3 Warranty expense for January 784 =11200*7%
4 Warranty expense for November 392
Warranty expense for December 1176
Less: Cost of replacements -630 =(14+28)*15
Estimated Warranty Liability account,
December 31, 2016
938
5 Estimated Warranty Liability account,
December 31, 2016
938
Warranty expense for January 784
Less: Cost of replacements -495 =33*15
Estimated Warranty Liability account,
January 31, 2017
1227